# Detecting Adversarial Data via Perturbation Forgery

### v2

@@ Introduction
Based on the theory support in the last section, we heuristically propose Perturbation Forgery to train a detector with strong generalization against all types of unseen attacks.
@@
As illustrated in Figure 2, Perturbation Forgery encompasses three elements: distribution perturbation, sparse mask generation, and pseudo-adversarial data production. At first, the noise distribution of a commonly used attack is continuously perturbed using Noise Distribution Perturbation into numerous noise distributions that remain proximal to the original, thereby forming the open covering. After that, we sample noises from the perturbed distribution and convert them into sparse local noise using Sparse Mask Generation. Subsequently, we convert half of the natural data into pseudo-adversarial data using Pseudo-Adversarial Data Generation and train a binary detector on the these data along with the remaining natural data.
Noise Distribution Perturbation
@@
Assuming the noise features form a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ (Du et al. 2022), we estimate the parameters mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ of $\mathcal{D}_{\mathrm{noise}}$ by
@@
where $\boldsymbol{m}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and $\boldsymbol{V}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d\times d})$ are vector and matrix of perturbation. $\alpha_{i}\sim\boldsymbol{U}(-\epsilon_{\mu},\epsilon_{\mu})$ and $\beta_{i}\sim\boldsymbol{U}(0,\epsilon_{\sigma})$ are scale parameters sampled from uniform distributions. After $n$ batches of training, we have a perturbed distribution set $\{\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})\}$, thereby forming the approximate open covering $\mathcal{P}_{c}$. Trained with $\mathcal{P}_{c}$, the detector can distinguish the distributions of natural and adversarial data.
Sparse Mask Generation
Unlike gradient-based adversarial attacks, generative-based attacks tend to generate imbalanced and anisotropy perturbations, showing much intensity in high-frequency and salient areas and mild in low-frequency and background areas. Besides, in the physical situation, adversarial attacks also tend to add patches to local areas. Noticing this, we propose sparse masks to perturbation forgery to assist the detector in detecting local adv-noises.
Previous research has demonstrated that the spectra of real and fake images differ significantly in high-frequency regions (Luo et al. 2021). Additionally, generative-based adversarial data manipulate characteristics, such as abnormal color aberrations, in these high-frequency regions (Wang et al. 2023a). To augment the detection performance on generative-based adversarial data, we need to sparse the pseudo-noise sampled from perturbed noise distribution and highlight the noise in high-frequency regions to mimic the style of generative-based attacks.
@@
Input: Training dataset, initial attack $\operatorname{Attack}$, saliency detection model $f_{s}$, and GradCAM model $f_{c}$.
Output: A detector with strong generalization capabilities against all types of unseen attacks.
1:  Generate adversarial noise using the initial attack $\operatorname{Attack}$ in Eq. 3.
@@
5:     for sample $\mathbf{x}_{i,k}$ half of batch $i$ do
@@
7:        Sample pseudo-adversarial noise $\widehat{\boldsymbol{\eta}}_{i,k}$ using Eq. 10.
8:        Generate pseudo-adversarial data $\widehat{\mathbf{x}}_{i,k}$ using Eq. 12.
@@
Given this, we first mask the high-frequency area of a natural sample using saliency detection (Qin et al. 2019) and GradCAM (Selvaraju et al. 2017):
$\operatorname{Mask}_{1}(\mathbf{x})=\operatorname{Map}(f_{s}(\mathbf{x});\gamma_{s})\cup\operatorname{Map}(f_{c}(\mathbf{x});\gamma_{c}),$ (6)
@@
$\operatorname{Map}(\mathbf{x};\gamma)=\operatorname{Clip}_{[0,1]}(\mathbf{x}-\gamma\mathbf{I}_{\mathcal{X}}).$ (7)
The noise that occurs in the high-frequency region tends to occupy only a small part of the region rather than spreading to the entire high-frequency spectrum. Therefore, we select the high-frequency point using the Sobel operator (Duda and Hart 1973) to sparsify the mask of sample $\mathbf{x}_{i,k}$:
$\operatorname{Mask}_{2}(\mathbf{x})=\operatorname{Map}(\operatorname{Sobel}(\mathbf{x});\gamma_{l}),$ (8)
@@
Pseudo-Adversarial Data Production
For the $i^{\mathrm{th}}$ batch of training, we sample pseudo-adversarial noise from the perturbed distribution by:
@@
where $h^{-1}:\mathbb{R}^{d}\rightarrow\mathcal{X}$ is an inverse function of $h$ to unflatten noise features, and $\gamma_{p}$ is a threshold that excludes data with a low likelihood. Note that $\boldsymbol{\eta}_{i}$ is a global noise and needs to be localized by a sparse mask. For natural data $\mathbf{x}_{i,k}$ in the $i^{\mathrm{th}}$ batch, the corresponding pseudo-adversarial data $\widehat{\boldsymbol{x}}_{i,k}$ is produced by adding the masked local noise to it:
@@
For each batch, we convert half of the natural data to pseudo-adversarial data by Perturbation Forgery and train a binary detector using the Cross-Entropy classification loss (Zhang and Sabuncu 2018). The overall training procedure with Perturbation Forgery is presented in Algorithm 1.
Experiments
To validate the detection performance of our approach against various adversarial attack methods, we conduct extensive experiments on multiple general and facial datasets.
Experiment Settings
Datasets and Network Architectures
The datasets for evaluation include 1) general dataset ImageNet100 (Deng et al. 2009), and 2) face datasets: Makeup (Gu et al. 2019), CelebA-HQ (Karras et al. 2017), and LFW (Gary et al. 2007). Gradient-based attacks and generative-based attacks are launched only on ImageNet100. For face attacks, we conduct Adv-Makeup on Makeup, AMT-GAN on CelebA-HQ, and physical attacks on LFW.
Adversarial Attacks
We adopt 9 gradient-based attacks, 5 generative-based attacks, and 6 face attacks to evaluate our method. For gradient-based attacks, BIM (Kurakin, Goodfellow, and Bengio 2016), PGD (Madry et al. 2017), RFGSM (Tramèr et al. 2018), MIM (Dong et al. 2018), DIM (Wu et al. 2021), NIM and SNIM (Lin et al. 2019), VNIM (Wang and He 2021), and Auto-Attack (Croce and Hein 2020) are used as the unseen attacks. For generative-based attacks, we adopt 3 GAN-based attacks including CDA (Naseer et al. 2019), TTP (Naseer et al. 2021), and M3D (Zhao et al. 2023), and 2 Diffusion-based attacks including Diff-Attack (Chen et al. 2023) and Diff-PGD (Xue et al. 2024) to craft adv-samples. For face attacks, 1 gradient-based attack TIPIM (Yang et al. 2021), 2 GAN-based attacks including AdvMakeup (Yin et al. 2021) and AMT-GAN (Hu et al. 2022), and 3 physical attacks including Adv-Sticker (Tong et al. 2021), Adv-Glasses (Sharif et al. 2016), and Adv-Mask (Tong et al. 2021) produce adversarial face samples. For each of the above attack methods, adversarial data is generated with the perturbation magnitude of $\epsilon=4/255$ (except for the unrestricted attacks) following (Zhang et al. 2023).
Table 2: Comparison of AUC scores (%) of detecting gradient-based attacks on ImageNet100.
Detector BIM PGD RFGSM DIM MIM NIM VNIM SNIM AA
LID (Ma et al. 2018) 0.9782 0.9750 0.9637 0.8942 0.9146 0.8977 0.8865 0.8752 0.9124
LiBRe (Deng et al. 2021) 0.9259 0.9548 0.9769 0.9243 0.8725 0.9013 0.8862 0.8694 0.8653
SPAD (Wang et al. 2023a) 0.9846 0.9851 0.9833 0.9815 0.9820 0.9823 0.9891 0.9877 0.9890
EPSAD (Zhang et al. 2023) 0.9998 0.9989 0.9992 0.9923 0.9918 0.9972 0.9946 0.9908 0.9998
ours 0.9911 0.9912 0.9911 0.9863 0.9931 0.9934 0.9878 0.9914 0.9941
Table 3: Comparison of AUC scores of detecting generative-based attacks on ImageNet100.
@@
CNN-Detection (Wang et al. 2020) 0.7051 0.6743 0.6917 0.3963 0.526
LGrad (Tan et al. 2023) 0.6144 0.6077 0.6068 0.5586 0.5835
Universal-Detector (Ojha, Li, and Lee 2023) 0.7945 0.8170 0.8312 0.9202 0.5531
DIRE (Wang et al. 2023b) 0.8976 0.9097 0.9129 0.9097 0.9143
@@
Baselines
We compare our method with state-of-the-art detection methods across three research fields, including adversarial detection, synthetic image detection, and face forgery detection, evaluating the detection of various types of adversarial attacks. For detecting gradient-based attacks, we compared our method with adversarial detection methods including LID (Ma et al. 2018), LiBRe (Deng et al. 2021), SPAD (Wang et al. 2023a), and EPSAD (Zhang et al. 2023). Since existing adversarial detection methods do not address generative-based adversarial data, we compared our method with synthetic image detection methods including CNN-Detection(Wang et al. 2020), LGrad(Tan et al. 2023), Universal-Detector(Ojha, Li, and Lee 2023), and DIRE(Wang et al. 2023b) for detecting generative-based attacks. For detecting GAN-based face attacks, we compared with face forgery detection methods including Luo et al. (Luo et al. 2021), He et al. (He et al. 2021), and ODIN (Liang, Li, and Srikant 2017). We conducted experiments using the pre-trained models and authors’ implementation code of the above methods.
Evaluation Metrics
Following previous adversarial detection works (Zhang et al. 2023), we employ the Area Under the Receiver Operating Characteristic Curve (AUC) as the primary evaluation metric to assess the performance of our classification models. The AUC is a widely recognized metric in the field of machine learning and is particularly useful for evaluating binary classification problems. For fair comparisons, the number of natural data and adversarial data in the experiments remained consistent.
Implementation details
XceptionNet (Chollet 2017) is selected as our backbone model and is trained on all the above datasets. For adv-noise distribution estimation, we utilize Torchattacks (Kim 2020) to generate 1000 noise images for each dataset. The uniform distribution parameters are set as $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$. The threshold of the sample control is set as $\gamma_{s}=0.05$. The thresholds of the binary indicator function are set as $\gamma_{s}=0$, $\gamma_{c}=125$ and $\gamma_{l}=100$. FGSM (Goodfellow, Shlens, and Szegedy 2014) is used as the initial attack for Perturbation Forgery. Training epochs are set to 10 and convergence is witnessed.
Detecting Gradient-based Attacks
We begin by comparing our method with state-of-the-art techniques in detecting gradient-based attacks on ImageNet100. All the attacks are seen to LID but unseen to other baselines and our method. For each attack, 1,000 adversarial images are generated in the white-box setting to test the detection models. All baseline methods are sourced from existing adversarial detection research. As reported in Table 2, although EPSAD performs well, our method demonstrates effective detection (achieving an AUC score of 0.9931 against MIM and 0.9914 against SNIM) across a variety of attack methods. These results validate our assumptions about the proximity of the noise distribution. By continuously perturbing the estimated noise distribution of a given attack (FGSM), we obtain an open covering noise of distributions from all the attacks. Given one seen attack, we are able to detect all the unseen attacks. The consistent detection performance verifies the effectiveness of our core idea of training a detector to distinguish the open covering of adv-noise distributions from natural ones and demonstrates the strong generalization capability of our method against gradient-based attacks. Additional results on CIFAR10 are provided in the supplementary materials.
Table 4: Comparison of AUC scores for detecting various adversarial attacks on the face.
@@
LID 0.6974 0.5103 0.6472 0.5319 He et al. (He et al. 2021) 0.5252 0.8823
LiBRe 0.5345 0.9739 0.8140 0.9993 Luo et al. (Luo et al. 2021) 0.6178 0.6511
EPSAD 0.9647 0.9462 0.9260 0.9578 ODIN(Liang, Li, and Srikant 2017) 0.6325 0.6970
@@
Detecting generative-based Attacks
To investigate the advantage of our method in detecting generative-based attacks, we compare with state-of-the-art methods of synthetic image detection and adversarial detection methods on ImageNet100. As shown in Table 3, our method substantially surpasses other methods, exhibiting highly effective detection capabilities. For adversarial detection baselines, due to the noise produced by generative models tending to be imbalanced compared to perturbations of traditional gradient-based attacks, these methods struggle to conduct effective detection. In contrast, with the assistance of sparse masks, our detector can perceive the local noises brought by GAN and diffusion models. As adversarial data is crafted from natural data, most synthetic image detection methods lose their effectiveness except DIRE(Wang et al. 2023b). This is probably because DIRE leverages the reconstruction error of images generated by diffusion models, thus maintaining a certain level of effectiveness.
Table 5: Ablation study on masks.

### v4

@@ 4 Perturbation Forgery
Based on the theory in the last section, we heuristically propose Perturbation Forgery to train a detector with strong generalization against various types of unseen attacks.
@@
As illustrated in Figure 2, Perturbation Forgery encompasses three elements: noise distribution perturbation, sparse mask generation, and pseudo-adversarial data production. At first, the noise distribution of a commonly used attack is continuously perturbed using noise distribution perturbation into numerous noise distributions that remain proximal to the original, thereby forming the open covering. After that, we sample noises from the perturbed distribution and convert them into sparse local noise using sparse mask generation. Subsequently, we convert half of the natural data into pseudo-adv-data using pseudo-adversarial data production. This pseudo-adv data replaces actual adversarial data and is combined with the remaining natural data to train a binary detector.
4.1 Noise Distribution Perturbation
@@
Assuming the noise features form a multivariate Gaussian distribution $p(\mathbf{z})=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ [10], we estimate the parameters mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$ of $\mathcal{D}_{\mathrm{noise}}$ by
@@
where $\boldsymbol{m}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and $\boldsymbol{V}_{i}\sim\mathcal{N}(0,\mathbf{I}_{d\times d})$ are vector and matrix of perturbation. $\alpha_{i}\sim\boldsymbol{U}(-\epsilon_{\mu},\epsilon_{\mu})$ and $\beta_{i}\sim\boldsymbol{U}(0,\epsilon_{\sigma})$ are scale parameters.After $n$ batches of training, we have a perturbed distribution set $\{\mathcal{N}(\widehat{\boldsymbol{\mu}}_{i},\widehat{\boldsymbol{\Sigma}}_{i})\}$, thereby forming the approximate open covering $\mathcal{P}_{c}$. Trained with $\mathcal{P}_{c}$, the detector can distinguish the distributions of natural and adversarial data.
4.2 Sparse Mask Generation
Unlike gradient-based adversarial attacks, generative-based attacks tend to generate imbalanced and anisotropic perturbations, showing much intensity in high-frequency and salient areas and mild in low-frequency and background areas. Besides, in physical scenarios, adversarial attacks also tend to add patches to local areas. Noticing this, we propose sparse masks to perturbation forgery to assist the detector in detecting local adv-noises.
Previous research has demonstrated that the spectra of real and fake images differ significantly in high-frequency regions [27]. Additionally, generative-based adversarial data manipulate characteristics, such as abnormal color aberrations, in these high-frequency regions [40]. To augment the detection performance on generative-based adversarial data, we need to sparse the pseudo-noise sampled from perturbed noise distribution and highlight the noise in high-frequency regions to mimic the style of generative-based attacks.
@@ 4.2 Sparse Mask Generation
Input: Training dataset, a commonly used attack, saliency detection model $f_{s}$, and GradCAM model $f_{c}$.
Output: A detector with strong generalization capabilities against various types of unseen attacks.
1:  Generate adversarial noise using the commonly used attack in Eq. 2.
@@
5:     for sample $\mathbf{x}_{i,k}$ in half of batch $i$ do
@@
7:        Sample pseudo-adv noise $\widehat{\boldsymbol{\eta}}_{i,k}$ using Eq. 9.
8:        Generate pseudo-adv data $\widehat{\mathbf{x}}_{i,k}$ using Eq. 11 and replace $\mathbf{x}_{i,k}$.
@@
Given this, for a natural sample $\mathbf{x}_{i,k}$ in the $\text{i}^{th}$ batch, we first mask the high-frequency area using saliency detection [33] and GradCAM [34]:
$\operatorname{Mask}_{1}(\mathbf{x}_{i,k})=\operatorname{Map}(f_{s}(\mathbf{x}_{i,k});\gamma_{s})\cup\operatorname{Map}(f_{c}(\mathbf{x}_{i,k});\gamma_{c}),$ (5)
@@
$\operatorname{Map}(\mathbf{x}_{i,k};\gamma)=\operatorname{Clip}_{[0,1]}(\mathbf{x}_{i,k}-\gamma\mathbf{I}_{\mathcal{X}}).$ (6)
The adv-noise that occurs in the high-frequency region tends to occupy only a small part of the region. Therefore, we select the high-frequency point from the gradient map obtained by Sobel operator [11] to sparsify the mask of sample $\mathbf{x}_{i,k}$:
$\operatorname{Mask}_{2}(\mathbf{x}_{i,k})=\operatorname{Map}(\operatorname{Sobel}(\mathbf{x}_{i,k});\gamma_{l}),$ (7)
@@
4.3 Pseudo-Adversarial Data Production
For the $i^{\mathrm{th}}$ batch of training, we sample pseudo-adv noise from the perturbed distribution by:
@@ 4.3 Pseudo-Adversarial Data Production
where $h^{-1}:\mathbb{R}^{d}\rightarrow\mathcal{X}$ is an inverse function of $h$ to unflatten noise features, and $\gamma_{p}$ is a threshold that excludes data with a low likelihood. Note that $\boldsymbol{\eta}_{i}$ is a global noise and needs to be localized by a sparse mask. The pseudo-adv data $\widehat{\mathbf{x}}_{i,k}$ is produced by adding the masked local noise to $\mathbf{x}_{i,k}$:
@@
For each batch, we convert half of the natural data to pseudo-adv data by Perturbation Forgery and train a binary detector using the Cross-Entropy classification loss [50]. The overall training procedure with Perturbation Forgery is presented in Algorithm 1.
Table 2: Comparison of AUROC scores of detecting gradient-based and adaptive attacks on ImageNet100 under $\epsilon=4/255$.
Detector BIM PGD DIM MIM NIM AA BPDA+EOT MM
LID [28] 0.9782 0.9750 0.8942 0.9146 0.8977 0.9124 0.9202 0.9227
LiBRe [8] 0.9259 0.9548 0.9243 0.8725 0.9013 0.8653 0.8714 0.8573
SPAD [40] 0.9846 0.9851 0.9815 0.9820 0.9823 0.9890 0.9885 0.9811
EPSAD [49] 0.9998 0.9989 0.9923 0.9918 0.9972 0.9998 0.9985 0.9923
ours 0.9911 0.9912 0.9863 0.9931 0.9934 0.9941 0.9927 0.9944
5 Experiments
To validate the detection performance of our approach against various adversarial attack methods, we conduct extensive experiments on multiple general and facial datasets. For additional experiments, analyses, and a discussion of limitations, please refer to the supplementary material.
5.1 Experiment Settings
Datasets
The datasets used for evaluation include: 1) general datasets: ImageNet100 [7] and CIFAR-10 [23] for evaluate gradient-based and generative-based attacks, and 2) face datasets: Makeup [16] for Adv-Makeup [46], CelebA-HQ [21] for AMT-GAN [19], and LFW [14] for TIPIM [45] and physical attacks.
Table 3: Comparison of AUROC scores of detecting generative-based attacks on ImageNet100.
@@ 5.1 Experiment Settings
      CNN-Detection [41]       0.7051       0.6743       0.6917       0.3963       0.5260
      LGrad [37]       0.6144       0.6077       0.6068       0.5586       0.5835
      Universal-Detector [32]       0.7945       0.8170       0.8312       0.9202       0.5531
      DIRE [42]       0.8976       0.9097       0.9129       0.9097       0.9143
      SPAD       0.9385       0.9064       0.8984       0.8862       0.8879
@@
Table 4: Comparison of AUROC scores for detecting various adversarial attacks on the face dataset.
@@
LID 0.6974 0.5103 0.6472 0.5319 He et al. [17] 0.5252 0.8823
LiBRe 0.5345 0.9739 0.8140 0.9993 Luo et al. [27] 0.6178 0.6511
EPSAD 0.9647 0.9462 0.9260 0.9578 ODIN[25] 0.6325 0.6970
@@
Adversarial Attacks
We adopt 5 gradient-based attacks, 3 adaptive attacks, 5 generative-based attacks, and 6 face attacks to evaluate our method. For gradient-based attacks, BIM [24], PGD [29], MIM [9], DIM [43], and NIM [26] are used as the unseen attacks. For adaptive attacks, we adopt AutoAttack (AA) [6], BPDA+EOT [18], and Minimum-Margin Attack (MM) [13]. For generative-based attacks, we adopt 3 GAN-based attacks including CDA [31], TTP [30], and M3D [51], and 2 Diffusion-based attacks including Diff-Attack [4] and Diff-PGD [44]. For face attacks, we adopt 1 gradient-based attack TIPIM [45], 2 GAN-based attacks including AdvMakeup [46] and AMT-GAN [19], and 3 physical attacks including Adv-Sticker [38], Adv-Glasses [35], and Adv-Mask [38].For each of the above attack methods, adversarial data is generated with the attack intensity of $\epsilon=4/255$ (except for the unrestricted attacks), and iterative attacks run for 10 steps using step size 2.
Baselines
To evaluate our method across various types of adversarial attacks, we compare it with state-of-the-art detection methods from three related research domains: adversarial detection, synthetic image detection, and face forgery detection. For gradient-based attacks, we benchmark against four adversarial detection methods: LID [28], LiBRe [8], SPAD [40], and EPSAD [49]. Since most adversarial detection methods are not designed to address generative-based adversarial attacks, we include a comparison with four synthetic image detection techniques: CNN-Detection [41], LGrad [37], Universal-Detector [32], and DIRE [42]. For detecting GAN-based face attacks, we also compare against two face forgery detection methods, Luo et al.[27] and He et al.[17], as well as ODIN [25].
Evaluation Metrics
Following previous adversarial detection works [49], we employ the Area Under the Receiver Operating Characteristic Curve (AUROC) [20] as the primary evaluation metric to assess the performance of our classification models. The AUROC is a widely recognized metric in the field of machine learning and is particularly useful for evaluating binary classification problems. For fair comparisons, the number of natural data and adversarial data in the experiments remained consistent.
Implementation details
XceptionNet [5] is selected as our backbone model and is trained on all the above datasets. For adversarial noise distribution estimation, we utilize Torchattacks [22] to generate 1000 noise images for each dataset. The uniform distribution parameters are set as $\epsilon_{\mu}=3$ and $\epsilon_{\sigma}=0.005$. The threshold of the sample control is set as $\gamma_{s}=0.05$. The thresholds of the binary indicator function are set as $\gamma_{s}=0$, $\gamma_{c}=125$ and $\gamma_{l}=100$. FGSM [15] is used as the initial attack, i.e. the commonly used attack, for Perturbation Forgery. Training epochs are set to 10 and convergence is witnessed.
5.2 Detecting Gradient-based Attacks
We begin by comparing our method with state-of-the-art techniques in detecting gradient-based attacks on ImageNet100. All the attacks are unseen by both the baseline methods and our method. For each attack, 1,000 adversarial images are generated in the gray-box setting, where the attacker has access to the victim model but is blind to the detection model. All baseline methods are sourced from existing adversarial detection research.
As reported in Table 2, although EPSAD performs well, our method demonstrates effective detection (achieving an AUROC score of 0.9931 against MIM and 0.9944 against MM) across a variety of attack methods. These results validate our assumptions about the proximity of the noise distribution. By continuously perturbing the estimated noise distribution of a given attack (FGSM), we obtain an open covering noise of distributions from all the attacks. Given one seen attack, we are able to detect all the unseen attacks. The consistent detection performance verifies the effectiveness of our core idea of training a detector to distinguish the open covering of adv-noise distributions from natural ones and demonstrates the strong generalization capability of our method against gradient-based attacks.
5.3 Detecting Generative-based Attacks
To investigate our method’s advantage in detecting generative-based attacks, we compared it with state-of-the-art synthetic image detection and adversarial detection methods on ImageNet100. As shown in Table 3, our method significantly outperforms the baseline methods, demonstrating highly effective detection capabilities.
For adversarial detection baselines, the imbalanced noise generated by GAN and diffusion models makes it difficult for these methods, which are typically designed for gradient-based attacks, to detect effectively. In contrast, with the help of sparse masks, our detector is able to detect the localized noise introduced by GANs and diffusion models.
Most synthetic image detection methods are less effective in detecting adversarial data crafted from natural samples, with the exception of DIRE[42]. This effectiveness may stem from DIRE’s use of reconstruction error from images generated by diffusion models, which maintains its robustness in this setting.
5.4 Detecting Face Adversarial Examples
Adversarial attacks targeting face recognition models represent an important branch of adversarial example research. Therefore, the detection ability of adversarial face examples is an important indicator of the performance of the detector. We perform various adversarial face attacks and compare our approach with state-of-the-art methods from adversarial face detection and face forgery detection.
As shown in Table 4, our method shows excellent performance against a variety of adversarial attacks targeting faces and surpasses many existing adversarial face detection techniques. It further indicates the ability of the sparse mask to capture the essential characteristics of face adversarial examples and the strong generalization capability of our method.
Table 5: Ablation study on masks (AUROC score), where $f_{s}$ denotes saliency detection model and $f_{c}$ represents the CAM model.